40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jul 5 20:57:04 2019, Koji, Update, SUS, Another M7.1 EQ Screenshot_from_2019-07-05_21-03-06.png
    Reply  Fri Jul 5 21:53:10 2019, Koji, Update, SUS, Another M7.1 EQ 
       Reply  Fri Jul 5 22:21:13 2019, Koji, Update, SUS, Another M7.1 EQ Screenshot_from_2019-07-05_22-19-57.png
          Reply  Fri Jul 12 07:37:48 2019, gautam, Update, SUS, M4.9 EQ in Ridgecrest 56.pngITMXunstick.pngITMX_UL.pdf
             Reply  Tue Jul 16 15:00:03 2019, gautam, Update, SUS, Multiple small EQs 
Message ID: 14729     Entry time: Fri Jul 5 22:21:13 2019     In reply to: 14728     Reply to this: 14755
Author: Koji 
Type: Update 
Category: SUS 
Subject: Another M7.1 EQ 

In fact, ETMX was not stuck until the M7.1 EQ today. After that it got stuck, but during the after shocks, all the OSEMs occasionally showed full swing of the light levels. So I believe the magnets are OK.

Attachment 1: Screenshot_from_2019-07-05_22-19-57.png  36 kB  Uploaded Fri Jul 5 23:22:37 2019  | Hide | Hide all
Screenshot_from_2019-07-05_22-19-57.png
ELOG V3.1.3-